Given: R = 1500.00 ft, Delta = 12 degrees 00'00" Right, and PC Station = 139+92.92. The azimuth from the PC to the PI was 40 degrees 00 minutes 00 seconds.

If a total station was setup at 142+00.00 on the curve, backsighting the PC, and 0 set, determine the angle right necessary to layout the PT of the curve.

Question 10 options:

183 degrees 57 minutes 18 seconds

192 degrees 00 minutes 00 seconds

186 degrees 00 minutes 00 seconds

3 degrees 57 minutes 18 seconds
Answer #1

- onu 533 R Civen, R: 1500, s. 12° pe station = 139 +92-92 Length of cowe, La 0.01745 33 RA i la 0.01745 33. (1500) (iz) - 3180 lection L2RR Deflection angle orenc Deflection angle 6 142700 = 180 ( 142700-1399292 (207.09 240 x 150 X (207.08) Po PC ePIDA sso 14270 13.955° PC 139 +92.92 6 92 ous - 22.073 -3.955. Tesco pt 143 +07.07 =3.0156W US Ica 143 07.08 Deflection ale to 142700 (107:08) -14200 QMA 180 2xTx1500 = 107.08 = Tom from PT = 2:045 figure Vom vetov